# Communication and Energy-Aware Multi-UAV Coverage Path Planning for Networked Operations

### v1

@@ III PROBLEM FORMULATION
Consider a ROI defined as a 2D polygonal area that needs to be fully scanned by a group of $N$ UAVs. Every point within the ROI must be observed at least once using the UAV’s sensor during the mission. The ROI may contain No-Fly Zones (NFZs), which are areas that UAVs must avoid. The primary objective of this UAV group is to work collaboratively to achieve complete coverage of the ROI while maintaining continuous connectivity within a UAV mesh network[31]. Each UAV must be able to communicate with all other UAVs at all times, either directly or through multi-hop communication.
Our goal is to generate multi-UAV coverage paths that ensure continuous inter-UAV communication, minimize energy consumption, maximize coverage efficiency, and avoid entering NFZs. The proposed communication and energy-aware mCPP algorithm is described at a high level in Section III-A. Section III-B explains the representation of the ROI on a grid, while Section III-C discusses the connectivity assessment in the coverage paths. The techniques used for estimating energy consumption are outlined in Section III-D, and Section III-E describes the iterative optimization process that identifies communication and energy-aware mCPP paths. For applications with short communication ranges, Section III-F proposes formation flying as a special case. Finally, Section III-G outlines the operational constraints and assumptions made during this work.
III-A Communication and energy aware mCPP Algorithm
Input: ROI, $N$, $\lambda$, $UAV_{param}$, $\mathcal{G}$, $\mathcal{O}$
Output: Best Paths ($P_{best}$)
1 $R^{*}\leftarrow+\infty$ // Best range requirement;
2 $E_{tot}^{*}\leftarrow+\infty$ // Toatal energy for $P_{best}$;
3 $f^{*}\leftarrow+\infty$ // Objective function;
4 $P_{best}\leftarrow\{\}$
5Process ROI to grid cells and NFZs;
6 Optimize grid alignment via shifts and rotations;
7 Optimization:
8       foreach trial in $N_{darp}$ do
9             Generate DARP initialisation keys;
10             Generate DARP paths;
11             Time-sample paths using $s_{f}^{u_{i}}$, $s_{t}^{u_{i}}$, and $r_{t}^{u_{i}}$;
12
13             Optimizing starting positions:
14                   foreach trial in $N_{shift}$ do
15                         Shift start position & compute $R_{req}$; Minimize $R_{req}$ via optimal shifts;
16                         Return optimal paths ($p$)
17
18
19            Compute $E_{tot}$ and $R_{req}^{{min}}$ for each set of paths;
20             Compute object. function $f_{o}=R_{req}^{{min}}+\lambda\cdot E_{tot}$;
21             if $f<f^{*}$ then
22                   Update $f^{*}\leftarrow f$;
23                   Update $R^{*}\leftarrow R_{req}^{{min}}$;
24                   Update $E_{tot}^{*}\leftarrow E_{tot}$;
25
26                  Update $P_{best}\leftarrow p$;
27
28            Update optimization model with results;
29
30
31
Algorithm 1 Communication and energy aware mCPP
The proposed communication and energy-aware mCPP algorithm generates flight paths for $N$ UAVs ($N\geq 2$) that maximize ROI coverage while maintaining continuous connectivity within the UAV mesh network. The method introduces a parameter $\lambda\in[0,\infty)$ to trade off between minimizing communication range and energy consumption. Setting $\lambda=0$ prioritizes minimizing communication range only, while higher values prioritize energy efficiency.
The algorithm requires inputs such as $UAV_{param}$: UAV parameters (e.g., turning radius, speed profile), $\mathcal{G}$: grid parameters (e.g., cell size, coverage threshold), and $\mathcal{O}$: optimization parameters (e.g., number of trials, optimization time). Algorithm 1 summarizes the process, which begins by discretizing the ROI into grid cells and incorporating NFZs. An optimal shift and rotation process adjusts the grid’s alignment and orientation to maximize ROI coverage, minimize energy consumption, and enable precise path planning [21].
The optimization process starts by generating DARP paths using random initialization on a 2D grid [9]. The objective function $f_{o}$ is evaluated for each path using (7), considering communication range and energy consumption.
The algorithm performs $N_{darp}$ optimization trials to identify effective DARP initializations, aiming to minimize $f_{o}$ for each path. The closed-loop structure of DARP paths allows for optimizing the mission starting point, further minimizing the communication range requirement.
A secondary optimization process with $N_{shift}$ trials adjusts UAVs’ starting positions to minimize the communication range ($R_{req}$) for maintaining connectivity. The minimum required range, $R_{req}^{{min}}$, is determined using (4).
After generating paths and identifying optimal starting positions, total energy consumption ($E_{tot}$) is estimated (Section III-D), and $f_{o}$ is computed using (7). If a trial yields a lower $f_{o}$, the optimal paths ($P_{best}$) and metrics are updated.
The iterative process continues for $N_{darp}$ trials, outputting UAV paths that balance communication reliability and energy efficiency based on the specified $\lambda$ value. Fig. 1 illustrates the optimization process.
III-B ROI and NFZ Representation on grid, Area division and Path planning
In mCPP, various methods have been proposed for area division and path planning. Conventional approaches, such as Trapezoidal and Boustrophedon decomposition, divide the ROI into sub-regions for each robot to cover. However, these methods have several limitations. They often result in an unbalanced distribution of coverage areas among the robots, which can lead to inefficient resource utilization. Moreover, these methods typically divide areas into simple polygonal shapes that not suited for maintaining connectivity during coverage, especially in arbitrary and complex ROIs.
To address these limitations, we propose using the DARP algorithm [9] for area division and path planning. DARP excels in dividing arbitrary and complex ROIs into intertwined sub-regions while maintaining a balanced area distribution among the robots. The intertwined area structure, combined with the iterative optimization of DARP initialization, helps generate sub-regions that are suited for connected coverage. The STC path planning algorithm[29] is then employed to generate closed-loop paths for covering the divided sub-regions.
ROI representation: The ROI is discretized into equal-sized square grid cells based on a specified path width. The cell side length is set to twice the parallel distance between paths, as required by the STC path generation algorithm . A coverage threshold $\tau\in[0,1]$ is applied to determine the inclusion of border cells in the ROI representation. Only cells meeting or exceeding this threshold are considered part of the ROI. For instance, $\tau=0.1$ includes border cells with at least 10% of their area inside the ROI, while $\tau=1.0$ requires cells to be entirely within the ROI, potentially reducing coverage efficiency. This same thresholding technique is applied to identify NFZs, ensuring accurate representation within the grid. To optimize coverage, we perform a shift and rotation operation that maximizes the number of cells meeting the coverage threshold, following the method detailed in [21]. The goal of this adjustment is twofold: (1) to reduce the number of partially covered cells at the edges of the ROI, thereby maximizing the usable coverage area, and (2) to minimize redundant overlaps, which can waste energy.
Area division and path planning: The DARP algorithm is employed for area division, and it utilizes the STC path planning algorithm, taking the discretized grid representation of the ROI as input. DARP divides the grid into exclusive sub-regions for each UAV to cover through a modified Voronoi partitioning, where grid cells are assigned to the nearest UAV based on initial positions. This process ensures the distribution of coverage areas among the UAVs according to the required ratio. The algorithm iteratively adjusts the boundaries between sub-regions to achieve the desired area distribution, resulting in a set of non-overlapping, connected sub-regions. Subsequently, the STC path planning algorithm is applied to generate closed loop coverage path for each sub region.
III-C Connectivity Assessment in Multi-UAV Paths
Time-sampling: The position of the $i^{th}$ UAV on the path at time $t$, denoted as $P_{t}^{u_{i}}$, is estimated through the time-sampling process shown in (1). This estimation is based on the speed parameters of each UAV, such as forward speed ($s_{f}^{u_{i}}$), turning speed ($s_{t}^{u_{i}}$), and turning radius ($r_{t}^{u_{i}}$). The individual time-sampling process allows each UAV to operate with distinct speed configurations, accommodating varying mission requirements and vehicle capabilities.
$P_{t}^{u_{i}}=f\left(s_{f}^{u_{i}},s_{t}^{u_{i}},r_{t}^{u_{i}}\right)$ (1)
Connectivity analysis: At any given time instance $t$, the required communication range $R_{req}(t)$ is the minimum distance $r$ that ensures all UAVs, with positions $P_{t}^{u_{i}}$, remain connected in a mesh network, as shown in (2). The overall range requirement $R_{req}$ for maintaining connectivity throughout the entire mission is then determined by finding the maximum communication range needed across all time instances $t\leq T$, where $T$ represents the total mission time, as expressed in (3).
$R_{req}(t)$ $=\min\{r:\{P_{t}^{u_{i}}\}\text{ form a mesh with range }r\}$ (2)
$R_{req}$ $=\max_{t\leq T}\left(R_{req}(t)\right)$ (3)
Optimizing launch point: The initial $R_{req}$ is calculated using the default starting point combination $s_{0}$. Due to the closed-loop nature of the generated paths, starting points can be shifted by $\Delta s_{i}$ in various combinations inside the path, each altering $R_{req}$. Finding the optimal starting point in this large set of combinations is computationally expensive and time-consuming. A probabilistic model-based optimization is particularly suited for this task, it efficiently explores the combination space by balancing exploration and exploitation. The high-level overview of this process is shown in Algorithm 2. Bayesian optimization is implemented over $N_{shift}$ trials to identify the starting point combination that minimizes $R_{req}$, as expressed in (4). This approach systematically finds the optimal starting points without exhaustively searching all combinations. The resulting $R_{req}^{{min}}$ is then used to calculate the objective function $f_{o}$.
$R_{req}^{{min}}=\min_{\{\Delta s_{i}\}}\left(\max_{t}\left(R_{req}(t,\{\Delta s_{i}\})\right)\right)$ (4)
Input: Time-Sampled Paths, $N_{shift}$, Step size ($s_{step}$)
Output: Best shift combination ($\{\Delta s_{i}\}_{best}$), Minimum required distance ($R_{req}^{{min}}$)
1 $\{\Delta s_{i}\}_{best}\leftarrow\{\};\ R_{req}^{{min}}\leftarrow+\infty$;
2 Optimization:
3       foreach trial in $N_{shift}$ do
4             Generate shifts $\{\Delta s_{i}\}$ and apply to paths;
5             $R_{req}\leftarrow\max_{t\in T}\left(R_{req}(t,\{\Delta s_{i}\})\right)$;
6             if $R_{req}<R_{req}^{{min}}$ then
7                   $R_{req}^{{min}}\leftarrow R_{req};\ \{\Delta s_{i}\}_{best}\leftarrow\{\Delta s_{i}\}$;

### v2

@@ III PROBLEM FORMULATION
We consider the problem of surveying a two-dimensional ROI $\mathcal{A}\subset{\mathbb{R}}^{2}$. The mission involves deploying a fleet of $N$ multirotor UAVs, each operating within the 2D plane above the ground. The position of the $i$-th UAV at time $t$ is denoted by ${\mathbf{x}}_{i}(t)\in\mathbb{R}^{2}$ and the set of positions of all UAVs over a horizon of $T$ seconds is referred to as a multi-UAV trajectory $\mathcal{T}:=\{{\mathbf{x}}_{i}(t)\mid 1\leq i\leq N,t\in[0,T]\}$. Each UAV is equipped with a sensor capable of observing an area on the ground, referred to as the sensor footprint and denoted by $\mathcal{F}({\mathbf{x}}_{i}(t))\subset{\mathbb{R}}^{2}$. The coverage requirement stipulates that every point within the ROI must be observed at least once by one of the UAV’s sensor during the mission, so that
$\bigcup_{i=1}^{N}\bigcup_{t\in[0,T]}\mathcal{F}({\mathbf{x}}_{i}(t))=\mathcal{A}.$ (1)
Notably, we do not require $\mathcal{A}$ to be convex or simply connected, and NFZs may exist within the convex hull of $\mathcal{A}$. The total energy consumed by all the $N$ UAVs when following a given multi-UAV trajectory $\mathcal{T}$ is denoted by $e(\mathcal{T})$. The UAVs are equipped with omnidirectional antennas and form a mesh network to maintain continuous connectivity [31]. Given $\mathcal{T}$, the communication radius required to ensure that all the UAVs can communicate with each other, either directly or through multi-hop paths involving other UAVs, is referred to as the connectivity radius and denoted by $r(\mathcal{T})$.
Solving the coverage problem will broadly require solving two sub-problems. Firstly, we want to generate a large number of collision-free trajectories that satisfy the coverage condition in (1). Secondly, among the trajectories that satisfy (1), we need to chose the trajectory $\mathcal{T}$ that minimizes $f_{o}(\mathcal{T}):=r(\mathcal{T})+\lambda e(\mathcal{T})$, where $\lambda$ is a weighting parameter that balances the relative importance of the required communication range and total energy consumption. In practice, however, we integrate these two phases, as the proposed algorithm will simultaneously generate and prune trajectories based on their objective function values.
Before describing the algorithm, we note that both $r(\mathcal{T})$ and $e(\mathcal{T})$ must be estimated for candidate trajectory $\mathcal{T}$ and the next two subsections describe the estimation routines used. While the resulting estimates are generally approximate—since they are oblivious to the environment—they are crucial because they enable us to systematically explore the space of valid trajectories.
III-A Estimation of total energy $e(\mathcal{T})$
There exist a number of approaches for estimating the energy consumption of UAV trajectories. For instance, the approach in [26] provides accurate energy estimates by utilizing detailed models for the UAV aerodynamics, motor, and battery. Although these detailed models can deliver precise energy estimates if available, they necessitate estimating a large number of UAV-specific parameters. Additionally, their high computational complexity makes it challenging to integrate them into the current path planning and optimization framework. Since UAVs maintain a constant speed for the majority of their flight path, we put forth a lightweight energy estimation algorithm that turns out to be sufficiently accurate.
The proposed energy estimation algorithm operates under two key modeling simplifications (a) the UAV paths consist solely of hovering, straight line segments covered at constant speed, and sharp turns, and (b) the UAV consumes constant powers $P_{h}$, $P_{f}$, and $P_{t}$ during hovering, forward flight, and turns, respectively. As we shall see later, the proposed algorithm will generate paths that adhere to the first assumption. The second simplification also introduces errors, e.g., due to wind, but the cumulative error is assumed to be negligible and ignored for the sake of simplicity. We further note that the more complicated energy estimation methods in [26, 10] also suffer from errors due to the presence of unknown winds.
Under these simplifications and for a given trajectory $\mathcal{T}_{i}$ of the $i$-th UAV, we simply need to measure the total time duration for which the UAV hovers ($T_{h}(\mathcal{T}_{i})$), moves along a straight line path ($T_{f}(\mathcal{T}_{i})$), and makes a sharp turn ($T_{t}(\mathcal{T}_{i})$) such that $T_{h}(\mathcal{T}_{i})+T_{f}(\mathcal{T}_{i})+T_{t}(\mathcal{T}_{i})=T$. The total energy consumption for trajectory $\mathcal{T}$ is given by
$e(\mathcal{T})=\sum_{i=1}^{N}P_{h}T_{h}(\mathcal{T}_{i})+P_{f}T_{f}(\mathcal{T}_{i})+P_{t}T_{t}(\mathcal{T}_{i}).$ (2)
The energy estimation model was validated through systematic real-world experimentation. Since the mission trajectories consist of only straight-line segments and turning maneuvers, we excluded the take-off and landing phases from our analysis. During validation, we recorded the UAV’s power consumption throughout the mission duration, enabling characterization of distinct power profiles for straight-line flight ($P_{f}$) and turning maneuvers ($P_{t}$). The total energy was then computed by integrating these profiles over their respective time intervals. While this simplified approach proves effective, its accuracy critically depends on two factors: the precision of power consumption calibration and the adherence of actual missions to our modeling assumptions. Notably, the model assumes constant-speed straight-line flight, symmetric turning maneuvers, and environmental conditions that closely match the calibration phase parameters. The model may require recalibration or exhibit reduced accuracy in scenarios involving variable flight speeds, asymmetric maneuvers, or significant environmental disturbances that deviate from the calibration conditions. Despite these constraints, the validation framework achieves a practical balance between computational efficiency and accuracy for our specific mission profiles. To verify the estimation model, we conducted three experiments on simple polygonal ROIs using a 3-UAV system. Test results showed total energy consumption of 134.44 Wh, 137.04 Wh, and 136.82 Wh, with individual UAV consumption ranging from 42.12-46.82 Wh per mission. Comparing with our estimated 130.44 Wh yielded an average accuracy of 95.8% across all tests, validating the model’s reliability for multi-UAV mission planning.
III-B Estimation of connectivity radius $r(\mathcal{T})$
We begin with formally defining $r(\mathcal{T})$ for a network of UAVs, represented by a time-varying weighted complete graph $\mathcal{G}(t)=(\mathcal{N},\mathcal{E},w_{t})$. The set of nodes $\mathcal{N}=\{1,2,\ldots,N\}$ represents the $N$ UAVs, and the edge set =$\mathcal{E}=\{(i,j)\mid i,j\in\mathcal{N},\,i\neq j\}$ includes all possible pairs of UAVs, reflecting potential communication links. The graph is dynamic due to its time-dependent weight function $w_{t}:\mathcal{E}\rightarrow\mathbb{R}_{+}$; specifically, the weight of edge $(i,j)$ at time $t$ is given by $w_{t}(i,j)=\|{\mathbf{x}}_{i}(t)-{\mathbf{x}}_{j}(t)\|$.
Overloading the notation, let $r(t)$ represent the minimum communication radius required to ensure that the mesh network formed by the UAVs is connected at time $t$. Equivalently,
$r(t)=\min\left\{r\geq 0\mid\text{the subgraph }\mathcal{G}_{r}(t)\text{ is connected}\right\}.$ (3)
where $\mathcal{G}_{r}(t):=(\mathcal{N},\mathcal{E}_{r}(t))$ is the subgraph induced by including only the edges with weights less than or equal to $r$, so that
$\mathcal{E}_{r}(t)=\left\{(i,j)\in\mathcal{E}\mid w_{t}(i,j)\leq r\right\}.$ (4)
In other words, $\mathcal{E}_{r}(t)$ consists of all edges that connect UAVs that are within $r$ of each other at time $t$. The subgraph $\mathcal{G}_{r}(t)$ is thus the graph where UAVs can communicate if they are within range $r$. Finally, the connectivity radius for the entire trajectory $\mathcal{T}$ is given by $r(\mathcal{T})=\max_{t\in[0,T]}r(t)$.
Given the graph $\mathcal{G}(t)$, we can determine $r(t)$ by identifying the minimum bottleneck spanning tree (MBST) of $\mathcal{G}(t)$. An MBST is a spanning tree in which the largest edge weight, known as the ’bottleneck’ edge, is minimized across all possible spanning trees of the graph. In general, for a complete graph, the MBST can be found using Camerini’s algorithm, which operates in $O(N^{2})$ time [32]. This approach is computationally feasible when $N$ is relatively small, typically manageable for $N$ below 10. Camerini’s algorithm identifies the MBST by iteratively selecting edges to form a spanning tree while minimizing the maximum edge weight, which aligns with the concept of reducing the ”bottleneck.”
However, when $N$ is large, an alternative method is required to maintain computational efficiency. In particular, $\mathcal{G}(t)$ is a Euclidean graph, as the weights $w_{t}(i,j)$ are derived from Euclidean distances. For Euclidean graphs, a minimum spanning tree (MST) can be efficiently computed in $O(N\log(N))$ time using Delaunay triangulation. Once the MST of $G(t)$ is computed, $r(t)$ can be found directly from the weight of the heaviest edge in the MST, which, by definition, is also the MBST.
To determine $r(\mathcal{T})$, we need to solve the global scalar optimization problem $\max_{t\in[0,T]}r(t)$. The grid-search algorithm, which finds $\max_{1\leq k\leq T/{\Delta t}}r(k\Delta t)$, is efficient for small $T$. However, for large $T$, we require a more efficient search algorithm that takes advantage of the special structure of the problem. To this end, we have the following result:
Theorem 1.
The communication radius $r(t)$ is $2v$-Lipschitz continuous, i.e., $|r(t)-r(t^{\prime})|\leq 2v|t-t^{\prime}|$ for any $t,t^{\prime}\in[0,T]$.
Proof:
We begin with introducing some notation. Let $\mathcal{G}_{\ell}(t)=(\mathcal{N},\mathcal{E}_{\ell}(t))$ denote the $\ell$-th spanning tree of $\mathcal{G}(t)$, where $\mathcal{N}$ is the set of UAVs (which remains constant over time), and $\mathcal{E}_{\ell}(t)$ is the set of edges in that spanning tree at time $t$. Let $r_{\ell}(t):=\max_{(i,j)\in\mathcal{E}_{\ell}(t)}\left\|{\mathbf{x}}_{i}(t)-{\mathbf{x}}_{j}(t)\right\|$ denote the weight of its bottleneck (longest) edge. By definition, $r(t)=\min_{\ell}r_{\ell}(t)$. Let $\ell^{\star}(t)=\arg\min_{\ell}r_{\ell}(t)$ denote the index of the MBST at time $t$, so $r(t)=r_{\ell^{\star}(t)}(t)$.
The proof begins with bound $r(t)-r(t^{\prime})$, which can be written as:
$r(t)-r(t^{\prime})$ $=r_{\ell^{\star}(t)}(t)-r_{\ell^{\star}(t^{\prime})}(t^{\prime})$ (5)
$=[r_{\ell^{\star}(t)}(t)-r_{\ell^{\star}(t^{\prime})}(t)]+[r_{\ell^{\star}(t^{\prime})}(t)-r_{\ell^{\star}(t^{\prime})}(t^{\prime})]$ (6)
$\leq[r_{\ell^{\star}(t^{\prime})}(t)-r_{\ell^{\star}(t^{\prime})}(t^{\prime})],$ (7)
where the inequality follows because $r_{\ell^{\star}(t)}(t)\leq r_{\ell^{\star}(t^{\prime})}(t)$, as $\ell^{\star}(t)$ is the MBST at time $t$ and therefore has the smallest bottleneck edge weight among all spanning trees at that time. Note that $\ell^{\star}(t^{\prime})$ is also a spanning tree but not necessarily the MBST at time $t$.
Next, consider the term $r_{\ell^{\star}(t^{\prime})}(t)-r_{\ell^{\star}(t^{\prime})}(t^{\prime})$. Since $\mathcal{E}_{\ell^{\star}(t^{\prime})}$ is the edge set of a specific spanning tree (the MBST at time $t^{\prime}$), and the set of UAVs $\mathcal{N}$ remains the same, the edge set $\mathcal{E}_{\ell^{\star}(t^{\prime})}$ does not change between times $t^{\prime}$ and $t$; only the positions of the nodes (and thus the edge weights) change. Therefore:
$r_{\ell^{\star}(t^{\prime})}(t)-r_{\ell^{\star}(t^{\prime})}(t^{\prime})$
$=\max_{(i,j)\in\mathcal{E}_{\ell^{\star}(t^{\prime})}}\left\|{\mathbf{x}}_{i}(t)-{\mathbf{x}}_{j}(t)\right\|-\max_{(i,j)\in\mathcal{E}_{\ell^{\star}(t^{\prime})}}\left\|{\mathbf{x}}_{i}(t^{\prime})-{\mathbf{x}}_{j}(t^{\prime})\right\|$
$\leq\max_{(i,j)\in\mathcal{E}_{\ell^{\star}(t^{\prime})}}\left|\|\mathbf{x}_{i}(t)-\mathbf{x}_{j}(t)\|-\|\mathbf{x}_{i}(t^{\prime})-\mathbf{x}_{j}(t^{\prime})\|\right|,$ (8)
where we used the fact that the difference between the maxima of two sets is less than or equal to the maximum absolute difference between corresponding elements, i.e., $\max_{k}a_{k}-\max_{k}b_{k}\leq\max_{k}|a_{k}-b_{k}|$. Note also that since the edge sets are the same at times $t$ and $t^{\prime}$, we can directly compare the edge weights.
It remains to bound the change in edge weights of a particular edge. Since each UAV moves at most $v\lvert t-t^{\prime}\rvert$ units over the time interval $\lvert t-t^{\prime}\rvert$, the change in the length of any edge $(i,j)$ is upper bounded by $2v\lvert t-t^{\prime}\rvert$. Hence we have been able to establish that $r(t)-r(t^{\prime})\leq 2v\lvert t-t^{\prime}\rvert$. Finally, by interchanging $t$ and $t^{\prime}$ in this argument, we obtain $r(t^{\prime})-r(t)\leq 2v\lvert t-t^{\prime}\rvert$. Combining both inequalities, we obtain the desired Lipschitz continuity result. ∎
Having established the Lipschitz continuity of $r(t)$, it is now possible to use the Shubert–Piyavskii method for finding its minimum. The Shubert–Piyavskii method is a one-dimensional minimization algorithm that can find the global minimum of a $2v$-Lipschitz continuous function up to an accuracy of $\epsilon$ within at most $\mathcal{O}(vT\epsilon^{-1})$ function evaluations, where the computational complexity of each iteration is fixed. Note that the grid search algorithm also achieves an accuracy of $\epsilon$ in $\mathcal{O}(T\epsilon^{-1})$ function evaluations, but in practice, for large $T$, the Shubert-Piyavskii method would generally be faster as it better exploits the structure of the function. However, if the time horizon $T$ is relatively small, the computational overhead associated with initializing and maintaining the approximations in the Shubert–Piyavskii method might not be justified, and the grid search algorithm may be competitive or even faster.
IV Proposed Approach
In this section, we detail the proposed approach of designing multi-UAV trajectories that minimize $f_{o}(\mathcal{T})$. We begin with describing the area division and path planning algorithms in Sec. IV-A. Next, Sec. IV-B describes the proposed optimization approach.
IV-A Area Division, and Path Planning
Given an ROI $\mathcal{A}\subset\mathbb{R}^{2}$ and the workload ratios $w_{i}$ (collected in the vector $\mathbf{w}$) of each of the $N$ UAVs such that $\sum_{i=1}^{N}w_{i}=1$, we consider the problem of generating $N$ coverage paths. This requires solving two coupled problems: (1) partitioning $\mathcal{A}$ into $N$ disjoint sub-regions, one for each UAV to cover, such that,
$\mathcal{A}=\bigcup_{i=1}^{N}\mathcal{A}_{i},\quad\mathcal{A}_{i}\cap\mathcal{A}_{j}=\emptyset\quad\forall i\neq j$ (9)
where $\text{Area}(\mathcal{A}_{i})/\text{Area}(\mathcal{A})=w_{i}$, and (2) generating trajectories $\{{\mathbf{x}}_{i}(t)\}_{i=1}^{N}$ that ensure complete coverage of each sub-region, forming the multi-UAV trajectory $\mathcal{T}$. Traditional decomposition methods such as trapezoidal and Boustrophedon are not suited for the partitioning step as the resulting partitions do not adhere to a specified workload distribution and are not designed to ensure inter-UAV connectivity as discussed in Sec. III-B. Instead, we utilize the DARP algorithm proposed in [9]. For the trajectory generation step, we will utilize the STC path planning algorithm [29].
To begin with, both DARP and STC algorithms require the ROI to be discretized into a uniform grid of square cells. Let us denote the discretized ROI as $\mathcal{R}$ containing cells indexed from 1 to $\lvert\mathcal{R}\rvert$. The side length of each cell is set to twice the sensor footprint $\mathcal{F}({\mathbf{x}}_{i}(t))$, as required by the STC path planning algorithm. The discretization process utilizes a tunable coverage threshold parameter, $\tau\in(0,1]$ which specifies the percentage of a cell’s area that must be inside $\mathcal{R}$ for it to be included. That is, a grid cell $j$ (which covers the square region $C_{j}$) is included in the discretized ROI $\mathcal{R}$ based on the following criteria:
$j\in\mathcal{R}\quad\text{if}\quad\frac{\text{Area}(C_{j}\cap\mathcal{A})}{\text{Area}(C_{j})}\geq\tau.$ (10)
For instance, if we set $\tau=1.0$, only fully covered cells may be included in $\mathcal{R}$, while lower values such as $\tau=0.1$ allow us to include cells that are at least 10% inside the ROI. The same thresholding technique applies to cells on the boundaries of the NFZs. To maximize coverage efficiency, we apply a shift and rotation operation to the grid following the optimization method in [21]. This operation minimizes partially covered cells at ROI edges, optimizing the usable coverage area.
IV-A1 Area Division
Given $N$ distinct seed cells $p_{1}$, $\ldots$, $p_{N}$ (collected in the vector $\mathbf{p}$), the DARP algorithm divides $\mathcal{R}$ into $N$ sub-regions using a modified Voronoi partitioning method. Let $\mathcal{R}_{i}$ denote the set of indices of the cells included in the $i$-th region. The DARP algorithm is initialized with $\mathcal{R}_{i}=\{p_{i}\}$ and iteratively includes a cell $c$ that is (a) a neighbor of the region $\mathcal{R}_{i}$, (b) not part of any other region, and (c) closest to $p_{i}$. Each region is updated in a round-robin fashion and the process continues until $|\mathcal{R}_{i}|/|\mathcal{R}|=w_{i}$ for all $i\in{1,\ldots,N}$. It is remarked that the DARP algorithm is not guaranteed to terminate for every $\mathbf{p}$. Further, the seed positions critically influence the shapes of $\{\mathcal{R}_{i}\}$ and hence the quality of the generated trajectories. We will therefore utilize an outer loop to find the optimal $\mathbf{p}$ that minimizes $f_{o}(\mathcal{T})$.
IV-A2 Path Planning
We utilize the STC path planning algorithm to generate closed-loop coverage paths within each $\mathcal{R}_{i}$. Together, the $N$ closed-loop paths form the multi-UAV trajectory $\mathcal{T}$. While the launch point of each UAV can be any point along the generated STC path, its position $k_{i}\in\mathcal{R}_{i}$ is also optimized to minimize $f_{o}(\mathcal{T})$. The launch points are collected into the vector ${\mathbf{k}}\in{\mathbb{R}}^{N}$.
In summary, a trajectory generated by the DARP and STC algorithms depends only on two $N\times 1$ parameter vectors, namely $\mathbf{p}$ and ${\mathbf{k}}$, and is denoted by $\mathcal{T}(\mathbf{p},{\mathbf{k}})$.
IV-B Optimal Trajectories
Finding the optimal trajectory amounts to solving
$\min_{\mathbf{p},{\mathbf{k}}}f_{o}(\mathcal{T}(\mathbf{p},{\mathbf{k}}))=r(\mathcal{T}(\mathbf{p},{\mathbf{k}}))+\lambda e(\mathcal{T}(\mathbf{p},{\mathbf{k}}))$ (11)
where the connectivity radius $r(\mathcal{T}(\mathbf{p},{\mathbf{k}}))$ is calculated as described in Sec. III-B and $e(\mathcal{T}({\mathbf{k}},\mathbf{p}))$ is calculated as described in Sec. III-A. Observe that there are $\binom{|\mathcal{R}|}{N}$ possible choices for initializing the DARP algorithm and $\prod_{i}|\mathcal{R}_{i}|$ possible launch points for the STC algorithm. Given the large and discrete search space, a joint optimization approach can be inefficient. Instead, we propose a hierarchical strategy that prunes large parts of the search space based on some heuristics.
The proposed algorithm proceeds in two nested loops. In the outer loop, we select a seed point $\mathbf{p}$ for DARP initialization and subsequently generate the paths using the STC algorithm. Observer that once the paths are generated, the energy $e(\mathcal{T})$ of the trajectory does not depend on the launch points. Hence, the inner loop goes through multiple UAV launch points ${\mathbf{k}}$ while seeking to find
${\mathbf{k}}(\mathbf{p})=\arg\min_{{\mathbf{k}}}r(\mathcal{T}(\mathbf{p},{\mathbf{k}}))$ (12)
where $r$ can be calculated as described in III-B. While generally there are $\prod_{i}|\mathcal{R}_{i}|$ possible launch points for ${\mathbf{k}}$, we utilize two key heuristics to reduce the search space. Firstly, grid search algorithm is utilized to find $\max_{1\leq i\leq T/\Delta t}r(i\Delta t)$. Secondly, all values of $r(t)$ that are calculated so far are maintained, and at any time if any $r(i\Delta t)$ exceeds the median of recorded values, the current inner iteration is skipped. Both optimization levels are implemented using the tree-structured Parzen estimator (TPE) [33] which is a Bayesian optimization framework. Fig. 1 illustrates this nested optimization process. In summary, the optimization process can be compactly written as
${\mathbf{k}}(\mathbf{p})$ $=\arg\min_{{\mathbf{k}}}r(\mathcal{T}(\mathbf{p},{\mathbf{k}}))$ (13)
$\mathbf{p}^{\star}$ $=\arg\min_{\mathbf{p}}f_{o}(\mathcal{T}(\mathbf{p},{\mathbf{k}}(\mathbf{p})))$ (14)
and is summarized in Algorithm 1.
Input: ROI $\mathcal{A}$, $N$, trade-off $\lambda$, UAV parameters
Output: Optimized multi-UAV trajectory $\mathcal{T}^{*}$
1
2Discretize $\mathcal{A}$ into grid $\mathcal{R}$;
3 Optimize grid alignment;
4 for $\ell=1$ to $N_{\text{darp}}$ do // Outer loop
5       Generate seed cells $\mathbf{p}$ via TPE;
6       Apply DARP($\mathbf{p}$) to partition $\mathcal{R}$ into $\{\mathcal{R}_{i}\}$;
7       Use STC on each $\mathcal{R}_{i}$ to form paths;